# Echoes of $R^{3}$ modification and Goldstone preheating in the CMB-BAO landscape

### v1

@@ Abstract
Introduction.— The amplitude of the scalar power spectrum $A_{s}$ and the spectral index $n_{s}$ in the baseline $\Lambda\mathrm{CDM}$ model provide powerful tests of inflationary dynamics. Combined with the tensor-to-scalar ratio $r$, the parameters $A_{s}$ and $n_{s}$ have ruled out several inflationary models using the Planck 2018 data [1]. The Starobinsky or the $R^{2}$ inflation [2, 3, 4], in which Einstein gravity is extended by an additional $R^{2}$ term, is one of the best-fit models to the Planck data. However, recent combined measurements (CMB-SPA+DESI) from the South Pole Telescope (SPT) [5], the Atacama Cosmology Telescope (ACT) [6], Planck and, baryon acoustic oscillation data from Dark Energy Spectroscopic Instrument (DESI) [7], found $n_{s}=0.9728\pm 0.0027$ [5]. This combined measurement is consistent with $n_{s}=0.9752\pm 0.0030$, found separately by the ACT collaboration using ACT, Planck (with lensing), and DESI-DR2 BAO data [6].
@@
The marginalised high value of $n_{s}$ excludes the $R^{2}$ model and the single-field regime of $R^{2}$-Higgs inflation at $2\sigma$ [5, 6]. In single-field attractor-type models, the predicted spectral index is $n_{s}=1-2/\mathcal{N}_{*}$, where $\mathcal{N}_{*}$ is the number of $e$-folds required for the reference scale (typically set at $k_{\rm ref}=0.05,\mathrm{Mpc}^{-1}$) to exit the horizon before the end of inflation. The parameter $\mathcal{N}_{*}$ depends explicitly on the post-inflationary reheating history, and with Standard Model (SM) field content, $\mathcal{N}_{*}$ in the $R^{2}$ and single-field-like regime of $R^{2}$-Higgs inflation lies within $[50,60]$ $e$-folds. This translates to $n_{s}\sim[0.9600\text{--}0.9667]$ and manifests as observed CMB-BAO tension [8, 9]. Several alternatives [10, 8, 11, 12, 13, 14, 15, 9, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] have been explored to reconcile the $R^{2}$ and $R^{2}$-Higgs models with the CMB-BAO tension, with the addition of an $R^{3}$ term being one of the simplest [32, 15, 33, 26]. In particular, it has been shown that the inclusion of an $R^{3}$ term can alleviate the tension in the $R^{2}$-Higgs model, especially when one deviates mildly from the single-field-like regime [33].
@@
In this follow up letter to Ref. [33], we show that Goldstone preheating in the $R^{3}$-modified $R^{2}$-Higgs inflation plays a quintessential role in explaining the observed high value of $n_{s}$. We demonstrate that Goldstone preheating (or equivalently, the production of longitudinal gauge bosons) induces parametric resonance leading to efficient production of Goldstone bosons. This results in significantly faster preheating than that of Higgs field alone. This rapid preheating improves the matching between inflationary and CMB scales and thereby helps alleviate the tension. We remark that all analyses of the $R^{2}$-Higgs model addressing the CMB-BAO tension thus far have adopted the unitary gauge, in which the Goldstone bosons are removed from the dynamics. We show that this gauge choice is ill-defined during every zero crossing of the Higgs background condensate after the end of inflation. By instead adopting a proper gauge choice, such as the Coulomb gauge, and employing a doubly covariant formalism for the scalar-field dynamics including all SM bosonic fields, we show that the gauge-invariant quantum energy densities of the produced Goldstone and Higgs quanta correctly reproduce the matching between the post-inflationary CMB reference scale $k_{\rm ref}$ and the inflationary scale.
The action and inflationary dynamics.— The action of $R^{3}$ modified $R^{2}$-Higgs inflation in Jordan frame is
@@
where $M_{\rm P}=\sqrt{1/\left(8\pi G\right)}\approx 2.4\times 10^{18}~\text{GeV}$ is the reduced Planck mass, $G$ is Newton’s constant and $\sqrt{-g_{J}}$ is the determinant of the mostly-plus metric. The $\Phi$ is the hypercharge +1+1 Higgs field, $\nabla_{\mu}=D_{\mu}+ig^{\prime}\frac{1}{2}Q_{Y}B_{\mu}+ig\,\bm{T}\cdot\bm{W}_{\mu}$ with $D_{\mu}$ is covariant space-time derivative. The $g^{\prime}$ and $g$ are the $U(1)_{Y}$ and $SU(2)_{L}$ couplings, $Q_{Y}$ is $U(1)_{Y}$ hypercharge and $\bm{T}$ are the weak isospin. Here, $\xi_{R}$ and $\xi_{c}$ are the dimension-four and -six self coupling of the Ricci scalar and $\xi_{H}$ is the nonminimal Higgs coupling. In the baseline $R^{2}$-Higgs model $\xi_{c}$ is zero.
@@
The $\phi$ and, the Higgs with decomposition $\Phi=(h+i\phi_{2},\phi_{3}+i\phi_{4})^{T}$, constitute a field-space manifold $\phi^{I}(x^{\mu})\in\{\phi,h,\phi_{2},\phi_{3},\phi_{4}\}$ with field-space metric components $G_{\phi\phi}=1$, $G_{hh}=e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{\rm P}}}$, $G_{\phi_{i}\phi_{i}}=e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{\rm P}}}$ with $i=2,3,4$. The presence of $G_{hh}$ and $G_{\phi_{i}\phi_{i}}$ is well known in the case of $R^{2}$-Higgs inflation and manifests itself as noncanonical kinetic terms [34, 35, 36, 37].
@@
The perturbations $\delta\phi^{I}$ are gauge dependent but the Mukhanov-Sasaki variables $Q^{I}=\delta\phi^{I}+(\dot{\varphi}^{I}/H)\psi$ are gauge independent, where $\psi$ is scalar perturbation from the metric [38, 39, 40]. The EoMs for $Q^{I}$ are [41, 35, 36, 37, 42, 43]
@@
where $\dot{\sigma}=\sqrt{G_{IJ}\dot{\varphi}^{I}\dot{\varphi}^{J}}$ and $\hat{\sigma}^{I}=\dot{\varphi}^{I}/\dot{\sigma}$.
We now consider a few benchmark points (BPs) in Table 1, mimicking a single-field-like regime with mild deviations [33], to examine whether they can yield a high $n_{s}$.
@@
The background EoMs Eq. (3) are solved with the initial conditions in Table 1. The EoMs of the perturbations in Eq. (4) on the other hand is solved in momentum space with Bunch-Davies (BD) initial conditions. The solutions are then inserted the into Eq. (5) to obtain $\mathcal{P}_{\mathcal{R}}$ and $n_{s}$. We find $\mathcal{P}_{\mathcal{R}}(k_{*})=2.128\times 10^{-9}$ $(2.104\times 10^{-9})$ and $n_{s_{*}}=0.9712$ $(0.9733)$ for BPaa (BPbb), with $k_{*}=3.0498\times 10^{-5}~M_{\rm P}$ $(1.3843\times 10^{-4}~M_{\rm P})$ and $\mathcal{N}_{*}=\ln a(t_{*})-\ln a_{\rm end}=-53.7$ (54.5)(-54.5), where $t_{*}$ corresponds to cosmic time when the reference mode $k_{*}$ exits horizon. These values lie within $\sim 1\sigma$ of the CMB-SPA+DESI constraints $\log(10^{10}A_{s})=3.0574\pm 0.0094$ and $n_{s}=0.9728\pm 0.0027$ [5]. The single-field estimate of $r\simeq 16\epsilon(\mathcal{N}_{*})$ yields $r\simeq 3.59\times 10^{-3}$ $(3.45\times 10^{-3})$, consistent with the 95% CL BICEP/Keck bound [44]. Matching $k_{*}$ to the reference scale $k_{\rm ref}$ is deferred for later section since it depends on the post-inflationary reheating history.
Preheating.— In the early part of reheating (i.e. preheating), the rapidly oscillating condensate can produce particles nonperturbatively and modify the thermal history, thereby affect in the matching. We show shortly that the BPs in Table 1 can induce successful preheating. To study preheating, a gauge choice is required. While the unitary gauge is commonly used in the $R^{2}$-Higgs model, it becomes ill-defined at each zero crossing of $h_{0}$ after inflation, as understood easily from Eq. (4) and from the explicit expression of $\mathcal{F}_{(\phi_{2})}$ [37, 42]. Same is true for $\phi_{3}$ and $\phi_{4}$. We therefore adopt the Coulomb gauge, $\partial_{i}Z^{i}=0$ and $\partial_{i}W^{\pm i}=0$, which is well defined for all background values [37, 42]. In this gauge one may treat either the Goldstone modes or the longitudinal gauge bosons as dynamical; in the following we keep the Goldstone modes dynamical and focus on $\phi_{2}$. The dynamics of $\phi_{3}$ and $\phi_{4}$ are analogous.
@@
where $m_{Z}^{2}=(g_{Z}^{2}/4)e^{-\sqrt{\frac{2}{3}}\frac{\phi_{0}}{m_{\rm P}}}h_{0}^{2}$, $M_{\phi_{2}}^{2}=m_{\mathrm{eff},(\phi_{2})}^{2}+m_{Z}^{2}$ and $(\prime)$ denotes the conformal time derivative. We have utilized here the Coulomb gauge condition $\partial_{i}Z^{i}=0$ and rescaled $Z_{0}\to Z_{0}/a$. The explicit form of $m_{\mathrm{eff},(I)}^{2}$, $\Upsilon$ and $\mathcal{K}_{Z}$ are provided in the appendix. The mode equation for the Higgs is found by setting terms with $\Upsilon$ to zero and the replacement $M_{\phi_{2}}^{2}\to m_{\mathrm{eff},(h)}^{2}$. The corresponding $M_{\phi_{2}}^{2}$ and $m_{\mathrm{eff},(h)}^{2}$ is plotted in Fig. 1. The magnitude of $M_{\phi_{2}}^{2}$ is much larger than $m_{\mathrm{eff},(h)}^{2}$ for respective BPs.
Figure 1: The evolution of $M_{\phi_{2}}^{2}$ and $m_{\mathrm{eff},(h)}^{2}$ for the BPs.
@@
We compute $\rho^{q}_{(\phi_{2})}$ by solving Eq. (7) with BD initial conditions (see appendix) in cosmic time via initializing all modes deep inside the horizon [37, 42]. Note that we solved Eq. (7) in cosmic time for computational convenience. The second term in the integrand of Eq. (8) is associated with the BD energy density for the corresponding mode. A mode is excited if $\rho_{k}^{(\phi_{2})}$ exceeds the corresponding BD energy density.
In Fig. 2, we show the quantum energy densities of the Higgs and Goldstone modes, $\rho^{q}_{(h)}$ (blue) and $\rho^{q}_{(\phi_{2})}$ (red), along with $\rho_{\rm inf}$ (black). For both BPs, $\phi_{2}$ triggers preheating first; at $\mathcal{N}_{\rm pre}\simeq 3.2$ and $1.4$ respectively. Higgs preheating is incomplete for BPaa, but completes for BPbb $\sim 2$ $e$-folds after inflation. Preheating is deemed complete if a perturbation energy density is equal to $\rho_{\rm inf}$ [42]. Note that the Goldstone preheating is faster due to larger $M_{\phi_{2}}^{2}$ and the $\Upsilon$-$\mathcal{K}_{Z}$ term in contrasts to only $m_{\mathrm{eff},(h)}^{2}$ in case of Higgs. The Goldstone preheating is faster for BPbb primarily due to larger $\xi_{H}$. Likewise, Higgs preheating is possible for BPbb due to its larger $m_{\mathrm{eff},(h)}^{2}$.
For the BPs, setting $\xi_{c}=0$ leads to incomplete preheating, reducing to pure $R^{2}$-Higgs inflation, which cannot account the observed high $n_{s}$. Preheating from $\phi_{3}$ and $\phi_{4}$ is nearly identical to $\phi_{2}$, differing only by $g_{Z}\to ie/(2\sqrt{2}s_{W})$. We find preheating from $\phi$ is weak, while transverse $Z$ and $W^{\pm}$ modes take longer. This finding is similar to Ref. [42]. For $\xi_{H}<1$, Goldstone as well as gauge boson preheating is subdued. In such scenario, nonvanishing $\xi_{c}$ can still explain high $n_{s}$, but additional fields are needed for preheating. Otherwise, thermalization proceeds via perturbative reheating [33, 23].
@@
Inflationary observables and scale matching.— We proceed now matching of $k_{*}$ (in $M_{\rm P}$) to the CMB reference scale $k_{\rm ref}/a_{0}=0.05~\rm{Mpc}^{-1}$, where $a_{0}$ is scale factor today. The CMB reference scale is defined as $k_{\rm ref}=k_{*}=a_{*}H_{*}$, giving [45]

### v2

@@ I Introduction
The amplitude of the scalar power spectrum $A_{s}$ and the spectral index $n_{s}$ in the baseline $\Lambda\mathrm{CDM}$ model provide powerful tests of inflationary dynamics. Combined with the tensor-to-scalar ratio $r$, the parameters $A_{s}$ and $n_{s}$ have ruled out several inflationary models using the Planck 2018 data [1]. The Starobinsky or the $R^{2}$ inflation [2, 3, 4], in which Einstein gravity is extended by an additional $R^{2}$ term, is one of the best-fit models to the Planck data. However, recent combined measurements (CMB-SPA+DESI) from the South Pole Telescope (SPT) [5], the Atacama Cosmology Telescope (ACT) [6], Planck and, baryon acoustic oscillation data from Dark Energy Spectroscopic Instrument (DESI) [7], found $n_{s}=0.9728\pm 0.0027$ [5]. This combined measurement is consistent with $n_{s}=0.9752\pm 0.0030$, found separately by the ACT collaboration using ACT, Planck (with lensing), and DESI-DR2 BAO data [6].
@@
The marginalised high value of $n_{s}$ excludes the $R^{2}$ model and the single-field regime of $R^{2}$-Higgs inflation at $2\sigma$ [5, 6]. In single-field attractor-type models, the predicted spectral index is $n_{s}=1-2/\mathcal{N}_{*}$, where $\mathcal{N}_{*}$ is the number of $e$-folds required for the reference scale (typically set at $k_{\rm ref}=0.05,\mathrm{Mpc}^{-1}$) to exit the horizon before the end of inflation. The parameter $\mathcal{N}_{*}$ depends explicitly on the post-inflationary reheating history, and with Standard Model (SM) field content, $\mathcal{N}_{*}$ in the $R^{2}$ and single-field-like regime of $R^{2}$-Higgs inflation lies within $[50,60]$ $e$-folds. This translates to $n_{s}\sim[0.9600\text{--}0.9667]$ and manifests as observed CMB-BAO tension [8, 9]. Several alternatives [8, 9, 11, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34] have been explored to reconcile the $R^{2}$ and $R^{2}$-Higgs models with the CMB-BAO tension, with the addition of an $R^{3}$ term being one of the simplest [35, 17, 36, 29]. In particular, it has been shown that the inclusion of an $R^{3}$ term can alleviate the tension in the $R^{2}$-Higgs model, especially when one deviates mildly from the single-field-like regime [36].
@@
In this letter, we show that Goldstone preheating in the $R^{3}$-modified $R^{2}$-Higgs inflation plays a quintessential role in explaining the observed high value of $n_{s}$. We demonstrate that Goldstone preheating (or equivalently, the production of longitudinal gauge bosons) induces parametric resonance leading to efficient production of Goldstone bosons. This results in significantly faster preheating than that of Higgs field alone. This rapid preheating improves the matching between inflationary and CMB scales and thereby helps alleviate the tension. We remark that all analyses of the $R^{2}$-Higgs model addressing the CMB-BAO tension thus far have adopted the unitary gauge, in which the Goldstone bosons are removed from the dynamics. We show that this gauge choice is ill-defined during every zero crossing of the Higgs background condensate after the end of inflation. By instead adopting a proper gauge choice, such as the Coulomb gauge, and employing a doubly covariant formalism for the scalar-field dynamics including all SM bosonic fields, we show that the gauge-invariant quantum energy densities of the produced Goldstone and Higgs quanta correctly reproduce the matching between the post-inflationary CMB reference scale $k_{\rm ref}$ and the inflationary scale.
II The action and inflationary dynamics
The action of $R^{3}$ modified $R^{2}$-Higgs inflation  [^1] We remark that in general $f(R)$ theory additional dimension-six operators, such as $\Phi^{2}R^{2}$ and $\Phi^{6}$, can also modify the inflationary predictions [37]. Moreover, additional dimension-four terms involving $R_{\mu\nu}R^{\mu\nu}$, $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$ and the Gauss-Bonnet term $R^{2}-4R_{\mu\nu}R^{\mu\nu}+R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$ are also possible. These terms along with their corresponding dimension-six operators may also modify inflationary predictions (See e.g. Refs. [38, 39, 40, 41, 42, 43, 44]). These effects are not considered here and being studied elsewhere. in Jordan frame is
@@ II The action and inflationary dynamics
where $M_{\rm P}=\sqrt{1/\left(8\pi G\right)}\approx 2.4\times 10^{18}~\text{GeV}$ is the reduced Planck mass, $G$ is Newton’s constant and $\sqrt{-g_{J}}$ is the determinant of the mostly-plus metric. The $\Phi$ is the hypercharge +1+1 Higgs field, $\nabla_{\mu}=D_{\mu}+ig^{\prime}\frac{1}{2}Q_{Y}B_{\mu}+ig\,\bm{T}\cdot\bm{W}_{\mu}$ with $D_{\mu}$ is covariant space-time derivative. The $g^{\prime}$ and $g$ are the $U(1)_{Y}$ and $SU(2)_{L}$ couplings, $Q_{Y}$ is $U(1)_{Y}$ hypercharge and $\bm{T}$ are the weak isospin. Here, $\xi_{R}$ and $\xi_{c}$ are the dimension-four and -six self coupling of the Ricci scalar and $\xi_{H}$ is the nonminimal Higgs coupling. In the baseline $R^{2}$-Higgs model the term associated with $\xi_{c}$ is zero.
@@
The $\phi$ and, the Higgs with decomposition $\Phi=(h+i\phi_{2},\phi_{3}+i\phi_{4})^{T}$, constitute a field-space manifold $\phi^{I}(x^{\mu})\in\{\phi,h,\phi_{2},\phi_{3},\phi_{4}\}$ with field-space metric components $G_{\phi\phi}=1$, $G_{hh}=e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{\rm P}}}$, $G_{\phi_{i}\phi_{i}}=e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{\rm P}}}$ with $i=2,3,4$. The presence of $G_{hh}$ and $G_{\phi_{i}\phi_{i}}$ is well known in the case of $R^{2}$-Higgs inflation and manifests itself as noncanonical kinetic terms [45, 46, 47, 48]. In the $\xi_{c}\to\infty$ limit $\tilde{\zeta}$ reduces to
$\tilde{\zeta}\simeq\xi_{R}+\frac{2}{\xi_{c}\,\xi_{R}}\left(\Theta-1-\frac{2\xi_{H}\,\Phi^{\dagger}\Phi}{M_{P}^{2}}\right).$ (3)
leading to to the baseline $R^{2}$-Higgs inflation potential as
$V_{E}(\phi^{I})=$ $e^{-2\sqrt{\frac{2}{3}}\frac{\phi}{M_{\rm P}}}\bigg[\frac{\lambda}{4}\left(h^{2}+\sum^{4}_{i=2}\phi_{i}^{2}\right)^{2}+\frac{M_{\rm P}^{4}}{4\xi_{R}}\Biggl\{1$-
$\qquad e^{\sqrt{\frac{2}{3}}\frac{\phi}{M_{\rm P}}}+\frac{\xi_{H}}{M_{\rm P}^{2}}\left(h^{2}+\sum^{4}_{i=2}\phi_{i}^{2}\right)\Biggr\}^{2}\bigg].$ (4)
@@
The perturbations $\delta\phi^{I}$ are gauge dependent but the Mukhanov-Sasaki variables $Q^{I}=\delta\phi^{I}+(\dot{\varphi}^{I}/H)\psi$ are gauge independent, where $\psi$ is scalar perturbation from the metric [49, 50, 51] as given as
$ds^{2}=$ $-(1+2\mathcal{A})dt^{2}+2a(t)(\partial_{i}\mathcal{B})dx^{i}dt+$
$a(t)^{2}\big[(1-2\psi)\delta_{ij}+2\partial_{i}\partial_{j}\mathcal{E}\big]dx^{i}dx^{j}.$ (6)
In the following we choose longitudinal gauge where the scalar perturbations $\mathcal{B}$ and $\mathcal{E}$ vanish. The EoMs for $Q^{I}$ are [52, 46, 47, 48, 53, 54]
@@
where $\dot{\sigma}=\sqrt{G_{IJ}\dot{\varphi}^{I}\dot{\varphi}^{J}}$, $\hat{\sigma}^{I}=\dot{\varphi}^{I}/\dot{\sigma}$ and $Q_{\sigma}=\hat{\sigma}_{I}Q^{I}$.
Figure 1: The $M_{\phi_{2}}^{2}$ (left) and $m_{\mathrm{eff},(h)}^{2}$ (right) and, their individual components as function of $\mathcal{N}$ for the BPaa (upper panel) and BPbb (lower panel).
We now consider a few benchmark points (BPs)  [^2] Note that, in general, a negative value of $\xi_{c}$ can reproduce the measured value of $n_{s}$, as discussed in Ref. [17, 29]. A detailed study of the case with negative $\xi_{c}$ is currently being carried out elsewhere. in Table 1, mimicking a single-field-like regime with mild deviations as in Ref. [36], to examine whether they can yield a high $n_{s}$.
@@
The background EoMs Eq. (5) are solved with the initial conditions in Table 1. The EoMs of the perturbations in Eq. (7) on the other hand is solved in momentum space with Bunch-Davies (BD) initial conditions. The solutions are then inserted the into Eq. (8) to obtain $\mathcal{P}_{\mathcal{R}}$ and $n_{s}$. We find $\mathcal{P}_{\mathcal{R}}(k_{*})=2.128\times 10^{-9}$ $(2.104\times 10^{-9})$ and $n_{s_{*}}=0.9712$ $(0.9733)$ for BPaa (BPbb), with $k_{*}=3.0498\times 10^{-5}~M_{\rm P}$ $(1.3843\times 10^{-4}~M_{\rm P})$ and $\mathcal{N}_{*}=\ln a(t_{*})-\ln a_{\rm end}=-53.7$ (54.5)(-54.5), where $t_{*}$ corresponds to cosmic time when the reference mode $k_{*}$ exits horizon. These values lie within $\sim 1\sigma$ of the CMB-SPA+DESI constraints $\log(10^{10}A_{s})=3.0574\pm 0.0094$ and $n_{s}=0.9728\pm 0.0027$ [5]. The single-field estimate of $r\simeq 16\epsilon(\mathcal{N}_{*})$ yields $r\simeq 3.59\times 10^{-3}$ $(3.45\times 10^{-3})$, consistent with the 95% CL BICEP/Keck bound [55]. Matching $k_{*}$ to the reference scale $k_{\rm ref}$ is deferred for later section since it depends on the post-inflationary reheating history. We note that one can define an effective single-field potential $W_{\rm eff}(\phi)$ by solving $\frac{\partial V_{E}}{\partial h_{0}}=0$ for $h_{0}$ and substituting back into $V_{E}$ to compute the background and perturbation EoMs. However, Ref. [36] shows that the resulting $n_{s}$ can differ significantly from that obtained using the full potential $V_{E}$. We also verified that isocurvature perturbations in this approach are about three orders of magnitude smaller for the reference mode and remain suppressed during inflation. We refer the reader to Ref. [36] for these details.
III Preheating
In the early part of reheating (i.e. preheating), the rapidly oscillating condensate can produce particles nonperturbatively and modify the thermal history, thereby affect in the matching. We show shortly that the BPs in Table 1 can induce successful preheating. To study preheating, a gauge choice is required. While the unitary gauge is commonly used in the $R^{2}$-Higgs model, it becomes ill-defined at each zero crossing of $h_{0}$ after inflation, as understood easily from Eq. (7) and from the explicit expression of $\mathcal{F}_{(\phi_{2})}$ [48, 53]. Same is true for $\phi_{3}$ and $\phi_{4}$. We therefore adopt the Coulomb gauge, $\partial_{i}Z^{i}=0$ and $\partial_{i}W^{\pm i}=0$, which is well defined for all background values [48, 53]. In this gauge one may treat either the Goldstone modes or the longitudinal gauge bosons as dynamical; in the following we keep the Goldstone modes dynamical and focus on $\phi_{2}$. The dynamics of $\phi_{3}$ and $\phi_{4}$ are analogous.
@@ III Preheating
where $m_{Z}^{2}=(g_{Z}^{2}/4)e^{-\sqrt{\frac{2}{3}}\frac{\phi_{0}}{m_{\rm P}}}h_{0}^{2}$, $M_{\phi_{2}}^{2}=m_{\mathrm{eff},(\phi_{2})}^{2}+m_{Z}^{2}$ and $(\prime)$ denotes the conformal time derivative. The $\mathcal{K}_{Z}$ and $\Upsilon$ are given as
The effective mass $m_{\mathrm{eff},(I)}^{2}$ is given as
$m_{\mathrm{eff},(I)}^{2}(\tau)=\mathcal{M}^{I}_{~~I}-\frac{1}{6}R_{E}G^{I}_{~I},$ (12)
where $R_{E}=-(2H^{2}+\dot{H})$ and
One can identify different components of $m_{\mathrm{eff},(I)}^{2}$ as
$m_{1,(I)}^{2}$ $=$ $G^{(I)J}(\mathcal{D}_{(I)}\mathcal{D}_{J}V_{E}),$ (14a)
$m_{2,(I)}^{2}$ $=$ $-\mathcal{R}^{(I)}_{\ \ JK(I)}\dot{\varphi}^{J}\dot{\varphi}^{K},$ (14b)
$m_{3,(I)}^{2}$ $=$ $-\frac{1}{M_{\rm P}^{2}a^{3}}\mathcal{D}_{t}\left(\frac{a^{3}}{H}\dot{\varphi}^{(I)}\dot{\varphi}_{(I)}\right),$ (14c)
$m_{4,(I)}^{2}$ $=$ $-\frac{R_{E}}{6},$ (14d)
with $(I)$ indices are not summed such that
$m_{\mathrm{eff},(I)}^{2}=\sum_{k}m_{k,(I)}^{2}.$ (15)
We have utilized here the Coulomb gauge condition $\partial_{i}Z^{i}=0$ and rescaled $Z_{0}\to Z_{0}/a$. The mode equation for the Higgs is found by setting terms with $\Upsilon$ to zero and the replacement $M_{\phi_{2}}^{2}\to m_{\mathrm{eff},(h)}^{2}$. The corresponding $M_{\phi_{2}}^{2}$ and $m_{\mathrm{eff},(h)}^{2}$ along with their respective contributions are plotted in Fig. 1. The dominant contribution to $M_{\phi_{2}}^{2}$ comes from $m_{Z}^{2}$, while $m_{1,(I)}^{2}$ dominates $m_{\mathrm{eff},(h)}^{2}$. Since both effective masses are evaluated at the background level and the Goldstone bosons are treated as perturbations, they receive no significant contribution from $V_{E}$. For the considered BPs, $M_{\phi_{2}}^{2}$ is much larger than $m_{\mathrm{eff},(h)}^{2}$.
@@
We compute $\rho^{q}_{(\phi_{2})}$ by solving Eq. (10) with BD initial conditions $s_{k}=e^{-ik\tau}/(\sqrt{2k\Delta_{(\phi_{2})}})$ and $s_{k}^{\prime}$. in cosmic time via initializing all modes deep inside the horizon [48, 53]. The $\Delta_{(\phi_{2})}$ factor in the denominator appears due to the presence of the second friction term in Eq. (10). Note that we solved Eq. (10) in cosmic time for computational convenience. The second term in the integrand of Eq. (16) is associated with the BD energy density for the corresponding mode. A mode is excited if $\rho_{k}^{(\phi_{2})}$ exceeds the corresponding BD energy density.
@@
In Fig. 2, we show the quantum energy densities of the Higgs and Goldstone modes, $\rho^{q}_{(h)}$ (blue) and $\rho^{q}_{(\phi_{2})}$ (red), along with $\rho_{\rm inf}$ (black). For both BPs, $\phi_{2}$ triggers preheating first; at $\mathcal{N}_{\rm pre}\simeq 3.2$ and $1.4$ respectively. Higgs preheating is incomplete for BPaa, but completes for BPbb $\sim 2$ $e$-folds after inflation. Preheating is deemed complete if a perturbation energy density is equal to $\rho_{\rm inf}$ [53]. Note that the Goldstone preheating is faster due to larger $M_{\phi_{2}}^{2}$ and the $\Upsilon$-$\mathcal{K}_{Z}$ term in contrasts to only $m_{\mathrm{eff},(h)}^{2}$ in case of Higgs. The Goldstone preheating is faster for BPbb primarily due to larger $\xi_{H}$. Likewise, Higgs preheating is possible for BPbb due to its larger $m_{\mathrm{eff},(h)}^{2}$. Here we take $\xi_{H}\sim 10$ as a representative value. However, larger values $\xi_{H}\gtrsim 10$ are also viable and, with successful preheating, may also assist in matching the scale.
In the absence of the $R^{3}$ term, the model reduces to the baseline $R^{2}$-Higgs potential of Eq. (4), as discussed above. While preheating may still occur for the ballpark values of $\xi_{R}$ and $\xi_{H}$ for the BPs [53] but, one cannot account the observed large $n_{s}$ in the absence $R^{3}$ term [36]. Preheating from $\phi_{3}$ and $\phi_{4}$ is nearly identical to $\phi_{2}$, differing only by $g_{Z}\to ie/(2\sqrt{2}s_{W})$. We find preheating from $\phi$ is weak, while transverse $Z$ and $W^{\pm}$ modes take longer. This finding is similar to Ref. [53]. For $\xi_{H}<1$, Goldstone as well as gauge boson preheating is subdued. In such scenario, nonvanishing $\xi_{c}$ can still explain high $n_{s}$, but additional fields are needed for preheating. Otherwise, thermalization proceeds via perturbative reheating [36, 25].
@@
IV Inflationary observables and scale matching
We proceed now matching of $k_{*}$ (in $M_{\rm P}$) to the CMB reference scale $k_{\rm ref}/a_{0}=0.05~\rm{Mpc}^{-1}$, where $a_{0}$ is scale factor today. The CMB reference scale is defined as $k_{\rm ref}=k_{*}=a_{*}H_{*}$, giving [56]